Status Report

NASA Space Station On-Orbit Status 5 February 2005

By SpaceRef Editor
February 5, 2005
Filed under , , ,
NASA Space Station On-Orbit Status 5 February 2005

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below. Saturday, first weekend rest day for the crew.

After wakeup at the regular 1:00am EST, FE Salizhan Sharipov again activated the PK-3/N turbopump of the Russian/German TEKh-20 Plasma Crystal-3 (PK-3) science payload in the Service Module (SM), to maintain the vacuum inside the ZB work chamber. The turbopump will be deactivated tonight at ~4:25pm before crew sleep.

Afterwards, Sharipov and CDR/SO Leroy Chiao performed the regular weekly 3-hr. task of thorough station cleaning. [“Uborka”, done every Saturday, includes removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of the Service Module (SM) dining table and other surfaces with “Fungistat” disinfectant and cleaning fan screens to avoid temperature rises.]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

The current run of EarthKAM (EK/Earth Knowledge Acquired by Middle School Students) came to an end today at 5:50am when the Science Officer powered off the payload in the Lab, then disassembled the camera equipment and returned it to stowage. The ground uplinked greetings and thanks from various schools for the highly successful run. [The EK program is operated by a staff of undergraduate students in the MOC (Mission Operations Center) at U. of Cal. at San Diego, using special software that they designed to make the flight information available to the middle school students and to process their image requests. During a mission, the MOC is staffed 8 hours/day to process requests, communicate with participating schools, and analyze and target the returned images. After the mission, staff members also annotate and create captions for images to highlight their educational applications for use in classrooms, slide shows, and for display. With students specializing in subjects ranging from Literature and Communications to Biology and Computer Engineering, the ISS EK Mission Operations team brings a unique array of talents to the program, making EK truly an outstanding feature of ISS utilization.]

For his “Saturday Science” program today, Dr. Chiao first reviewed procedures for the DAFT (Dust and Aerosol Measurement Feasibility Test) investigation, then tagged up with ground ops personnel to discuss the project, which is concerned with measuring dust concentration in the air. Afterwards, Leroy loaded the experiment’s “P-Trak” software on the EXPRESS Rack laptop. [DAFT’s P-Trak and DustTrak are diagnostic devices assisting researchers in the development of SAME, the Smoke Aerosol Measurement Experiment. Its parent hardware has an opportunity to fly on ISS mission 13A.1 if the DAFT-1 and DAFT-2 tests are successful.]

Leroy also completed his daily check of the Total Dose reading and End File values of the TEPC (Tissue Equivalent Proportional Counter), which he relocated to the Node on 1/31, and called the data down at the evening DPC (daily planning conference). [This is currently a daily requirement since the UOP (utility outlet panel) near the TEPCs temporary location (for two weeks) does not allow automated telemetry monitoring by the ground.]

The FE completed the routine SOZh/ECLSS servicing/inspection in the SM, including the ASU toilet facilities.

The crew performed their daily 2.5-hr. physical exercise program on TVIS treadmill (aerobic), RED exerciser (anaerobic) and VELO cycle with bungee cord load trainer (both aerobic and anaerobic). Salizhan’s daily protocol currently prescribes a four-day microcycle exercise with 1.5 hr on the TVIS (today: Day 2 of a new set) and one hour on VELO.

Chiao then transferred the daily TVIS and RED exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data, followed by their erasure on the HRM storage medium.

At ~5:30am EST, the crew held their weekly teleconference with ISS Program Management at JSC/Houston via S-band/audio.

Later, at ~8:55am, Chiao and Sharipov engaged in their weekly planning conference (WPC) with the ground, discussing next week’s “Look-Ahead Plan” (prepared jointly by MCC-H and TsUP/Moscow planners), via S-band/audio, reviewing upcoming activities and any concerns about future on-orbit events.

Working off the Russian task list, Sharipov was to conduct, at his leisure, another one of the regular monthly sessions of the VC6 “Delta” program’s ETD experiment (Investigation of the Coordination of Eye and Head Movements). [After a calibration with the calibrating unit, the experiment investigates horizontal eye and head movement coordination, measured Listing’s plane, and determined the orientation of the vestibulo-ocular coordinate system, using five target marks on the horizontal plane. Each step required another prior calibration run, using visual target cues or the calibration unit.]

Also added to his task list was Salizhan’s first run of the Russian Ekon earth photography program for recording environmental damages with the Nikon D1 (800mm lens) camera, for which he had performed a hardware functionality check with laptop TP2 on 12/27. [EKON (KPT-3) is an experiment by the Russian Environmental Protection Services, which uses the Nikon D1 camera for observation and imaging of selected ground targets such as oil spills. Today’s targets were the Pacific Ocean, the Red Sea and the Persian Gulf.]

A third new item in Sharipov’s “job jar” was the long-term recurring task of imaging the externally mounted PKZ-1V Kromka 1-3 contamination experiment tablet. [The Kromka tablet, deployed on handrail 2614 of the DC-1 “Pirs” docking compartment, collects thruster plume effluents. The pictures are taken with the Kodak 760 digital still camera (DSC) from the EVA hatch 1 “illyuminator” (window) in the DC-1.]

As of last night (5:00pm), the Elektron O2 generator was functioning nominally in 50 amps mode, after both its primary and backup micropumps had shut down yesterday three times over 24 hours. [After the first two times, the crew was instructed to perform the restart procedure, take the Elektron to the 50-amp mode, wait 30 minutes and then take it back to the 32-amp mode. The machine would stay on for several hours in the 32-amp mode and then the pumps would shut down again. The third time, TsUP/Moscow instructed the crew to restart the unit in 50-amp mode and keep it there overnight.]

A new Shuttle prepack list has been uplinked. As reported by the CDR last night, cargo prepacking for the LF-1 Shuttle mission (STS-114), the first RTF (return-to-flight) checkout flight, has made good progress. As of today, 65% of the items approved to be prepacked have been completed. The crew will continue packing through the end of the Increment.

The dry cargo volume of Progress M-51 (16P) has been completely unloaded and is ready to be reloaded with items to be disposed. The crew had a tagup with the specialists to confirm the IMS (Inventory Management System) updates related to the completion of the unloading.

There are currently nine ThinkPad A31p NGLs (Next Generation Laptops) and one ThinkPad 760XD (the OCA Router) in use aboard the station. Troubleshooting of one additional A31p shell (#1002, SSC8) is planned for the near future. Five 760XD laptop shells are stowed, including three failed.

A conjunction (close encounter) with the Explorer 8 satellite (launched November 1960) occurred this morning at 6:59am. There was no probability of collision (PC), and no avoidance maneuver was required.

A station reboost is planned for 2/15 (~8:22pm EST), for about 1.8 m/s delta-V, to set up orbital phasing for the Progress 17 launch on 2/28.

Weekly Science Update (Expedition Ten — 14th):

GASMAP: Nothing new.

Human Research Facility/Workstation (HRF WS): Continuing.

Advanced Ultrasound (ADUM): A new ADUM session is planned for next week. FE Sharipov will start it off with an OPE session, followed by two scan sets later in the week. The PI and experiment teams are looking forward to another great week of ultrasound activities on the ISS.

Hand Posture Analyzer (HPA: Nothing new.

In-Space Soldering Investigation (ISSI): Operations are complete.

Pore Formation and Mobility Investigation (PFMI): Nothing new.

Space Acceleration Measurement System (SAMS): Nothing new.

Microgravity Acceleration Measurement System (MAMS): Nothing new.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES): PCG-STES is performing nominally.

Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS): Nothing new.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE): Nothing new.

Binary Colloidal Alloy Test-3 (BCAT-3): The Science Officer will perform another photography session for BCAT-3 on 2/8. The team is looking forward to seeing the photographs of Samples 4, 5, 8, 9 and 10.

Renal Stone (RS): Nothing new.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES): Nothing new.

Foot/Ground Reaction Forces during Space Flight (FOOT): Nothing new.

Materials ISS Experiment (MISSE): In progress. Deployed outside on the U.S. Airlock. Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI): The CBOSS PI and team thanked Dr. Chiao again for choosing FDI for his “Saturday Science” activity last week. It proved to be an effective method at removing bubbles from the TCM, and the team is currently working to have a procedure that documents the bubble removal technique through pictures. This technique will also be incorporated into future cell culture experiments.

Dust and Aerosol Measurement Feasibility Test (DAFT): Leroy was thanked for choosing DAFT for today’s “Saturday Science” activity. The team was looking forward to talking with Chiao during today’s crew conference. “Hopefully we will have the opportunity to obtain some data points soon”.

Biotechnology Specimen Temperature Controller (BSTC): Nothing new.

Yeast Group Activation Packs (Yeast GAP): Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM): Initial setup and activation of the ISS EarthKAM 760 digital camera occurred on 2/1 (3:25am EST). The first test image was successfully commanded at 5:28am that day.. The pictures received on the ground using the 180mm lens were clear and wonderfully detailed. 87 of 137 participating middle schools from across the U.S., Spain, Italy, Mexico, Canada, Germany, Puerto Rico, and Japan have successfully requested images. School participation ended last night at 6:00pm.

Earth Science Toward Exploration Research (ESTER): Nothing new.

Serial Network Flow Monitor (SNFM): Nothing new.

Fluid Merging Viscosity Measurement (FMVM): Nothing new.

Space Experiment Module (SEM): The SEM team thanked the crew again for last week’s video taping. This introduction to SEM “went off without a hitch, and the entire team is grateful for your effort”. Participants are looking forward to the next set of photos.

Viscous Liquid Foam–Bulk Metallic Glass (Foam): Nothing new.

Effects of Prolonged Space Flight on Human Skeletal Muscle (BIOPSY): Nothing new.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2): Planned.

Commercial Generic Bioprocessing Apparatus (CGBA): Nothing new.

Miscible Fluids in Microgravity (MFMG): Nothing new.

Educational Payload Operations (EPO): Nothing new.

Capillary Flow Experiment (CFE): Nothing new.

Crew Earth Observations (CEO): The ground has to date received a total of 8853 CEO images. The bulk of recent imagery documents the crew’s encounters with the wintry, poorly illuminated Northern Hemisphere. Nevertheless investigators are finding a number of striking views of cities in snow as well many frozen or partially frozen water bodies. The views of the snowbound northeastern US with the long lens are especially good. Elsewhere, a recent view of the Issaouane Erg (sand sea) located in eastern Algeria will be posted on NASA’s Earth Observatory website this weekend. With fantastic detail, this low-light image illustrates the variety of dune formations and types in the very active area of the Erg. The crew’s “imagery and technique continue to improve”, and they are “patiently acquiring many of our targets as weather and lighting permits”.

Today’s CEO (crew earth observations) photo targets, no longer limited in the current LVLH attitude, were North Mariana Islands, & Guam (weather was predicted to be clear for photography of reefs in the southern islands of the Mariana chain. The island of Guam should also have been visible to the south of Rota. Detailed nadir images of the islands and reefs are useful for production of accurate maps and assessment of reef health), Manila, Philippines (this overpass provided an opportunity for detailed photography of this island megacity. Overlapping swaths of the urban core are useful to track growth, land cover, and land use change. This information provides necessary inputs into urban climate models), Dhaka, Bangladesh (this nadir overpass provided an opportunity for high resolution mapping of this urban center. Overlapping mapping swaths along the urban-rural fringe are useful for assessment of potential geohazards resulting from flooding), and Irrawaddy River Delta, Burma (this overpass brought the ISS over the upper reaches of the Irrawaddy delta. Mapping swaths along the rivers draining into the delta are useful for monitoring of hydrologic change resulting from land use changes in the watershed. These data are also useful for prediction of potential flooding hazards).

CEO photography can be viewed and studied at the websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 10 crew visit:

Expedition 10 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Upcoming Key Events:

  • ISS Reboost — 2/15 (~8:22pm EST, ~1.8 m/s; phasing for 17P launch);
  • Progress 16 (16P) undocking & destructive reentry — 2/27;
  • Progress 17 (17P) launch — 2/28;
  • EVA-13 — 3/25;
  • Soyuz 10 (10S) launch — 4/15 with Expedition 11 (CDR Sergei Krikalev, FE/SO John Phillips);
  • Soyuz 9 (9S) undock — 4/25 with Exp. 10 crew (after 193 days on orbit, 191 days on board ISS);
  • Progress 18 (18P) launch — 6/10;
  • Progress 19 (19P) launch — 8/24;
  • Soyuz 11 (11S) launch — 9/27.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit (as of this morning, 6:11am EST [= epoch]):

  • Mean altitude — 357.2 km
  • Apogee height — 363.7 km
  • Perigee height — 350.7 km
  • Period — 91.69 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0009694
  • Solar Beta Angle — -63.9 deg (magnitude decreasing)
  • Orbits per 24-hr. day — 15.70
  • Mean altitude loss in last 24 hours — 60 m
  • Revolutions since FGB/Zarya launch (Nov. 98) — 35497

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.